Consider an automobile moving at v mph that skids d feet after its brakes lock. Calculate how far it would skid if it was moving at 2v and the brakes were locked.   







1
Expert's answer
2021-12-09T09:22:35-0500

By conservation of energy:


"\frac{mv^2}2=\mu mgd,\\\space\\
\frac{m(2v)^2}2=4\frac{mv^2}2=\mu mg(4d)."

4 times further.